OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 5x = x2 - 24? HELP?

OpenStudy (zpupster):

you posted the same question twice

OpenStudy (anonymous):

i never got an answer

OpenStudy (zpupster):

change the equation --set y on one side can you do that??

OpenStudy (anonymous):

yes y=x2-24-5x

OpenStudy (zpupster):

or y= x2 - 5x -24 is a little better now can you factor the right side?

OpenStudy (anonymous):

i dont know how

OpenStudy (zpupster):

do you know the Foil method

OpenStudy (anonymous):

no

OpenStudy (zpupster):

ok usually you learn that first but no problem we will walk thru it

OpenStudy (zpupster):

y= x2 - 5x -24 here is the binomial factors for the right hand side (x-8)(x+3) now when we use the FOIL method mult ---first,outside,inside,last first=x2 outside=3x inside=-8x last = -24 combine outside and inside terms = -5x y= x2 - 5x -24 (x-8)(x+3) =0 we set y =0 now there is a property-- I forget but you set each of the binomials to zero x-8=0 x+3=0 you can solve for x and get the x intercepts understand?? this is the same process you can use for some of your other questions you posted